# Labelling Data with Unknown References

### v3

@@ Appendix A Detailed Proofs
In this section we provide more careful treatments of the proofs of Lemma 1 (Appendix A.1) and Theorem 1 (Appendix A.4).
@@ A.1 Correctness of the EV protocol
We separate our proof in two: the first part concerns itself with correctness, or whether the challenges will provide sufficient information to determine if the evaluator can be lying. The second part deals with robustness under lying, cheating, or lucky scenarios by the evaluator.
A.1.1 Part 1 (Correctness)
@@
Lemma 2.
Let $x,x^{\prime}\in X$ be two binary strings such that $x\neq x^{\prime}$. In order to test whether an evaluator knows $f(x)=f(x^{\prime})$ without calling $f$, it is necessary to challenge the evaluator in both $x\equiv x^{\prime}$ and $x\cong x^{\prime}$. Moreover, these conditions are sufficient.
@@
(Necessary) It follows from the definition of $C$. Since $f(x)=f(x^{\prime})$, testing $x\equiv x^{\prime}$ is a necessary condition, but not sufficient. This challenge yields information around the internal structure of $x,x^{\prime}$ up to permutation. It is necessary because it is possible to have two $x,x^{\prime}$ such that $x^{\prime}\not\in\{\pi_{i}(x)\}$ (i.e., two strings that aren’t permutations of one another), and yet $C(x^{\prime})=C(x)$. However, this does not prove that x is similar to x’, since there could be a subset $t\in S_{x^{\prime}}.t\not\in S_{x}$ and $c(t)=c(s)$ for some $c\in\mathcal{C}$ (for example, if $c$ contains a disjunctive clause). Hence running this challenge provides information as to whether the encoding of $C(x^{\prime})$ contains all information fed into $\sigma$ for a specific $x,x^{\prime}$. It accounts for variation in internal structure of the string so that $C(x)=C(x^{\prime})$, but does not account for the ordering of these bitstrings.
@@
(Sufficient) It follows directly from the earlier statement. Suppose there exists a pair $x,x^{\prime}$ such that either (or both) $\not\equiv x^{\prime}, x≇x$$\not\cong x^{\prime}$ and also $f(x)=f(x^{\prime})$. This would imply that (1) either the string , a contradiction.
@@
While a natural consequence of Lemma 2 is that only checking for isomorphism is required for testing equality, we note that the goal of the challenges is to determine if the evaluator is lying, not the label itself.
@@ A.1.2 Part 2 (Robustness)
There are then two (disjoint) possibilities for a generated $\langle x^{\prime},y^{\prime}\rangle$: either that the evaluator is lying, but gets lucky; or that it is cheating and can generate the answer for one of the challenges. It is worth noting that generating an answer to pass both challenges is equivalent to knowing $f$.
@@ A.3 Lie 2: $f(x^{\prime})\neq y^{\prime}$
It follows from Lemma 2 that we know that both challenges together are sufficient to ascertain the truthfulness of a label; and from Lemma 3 we know that the ability of the verifier to do this is bounded. Putting both lemmas together we obtain the proof for Lemma 1.
@@ A.4 Correctness Bounds of the No-Data Algorithm
For this, it is easier to work with errors $\epsilon=1-\alpha$, and remark that, from Lemma 1 the first term (undetected, no flip) is given by
@@ Appendix B Characterisation Bounds
Straightforward. ∎
@@ Appendix C Extensions
C.1 Distance-based Constraints
The constraint that $C(x)=C(x^{\prime})$ is too strict for some applications. We sketch out in this section a variant of the No-Data Algorithm where, instead of requiring equality, we request closeness with respect to some arbitrary metric $\delta$ in Challenge 2. Suppose now that the failure is given whenever $\delta(C(x),C(x^{\prime}))\geq\epsilon$, for some $\epsilon\in[0,1/2]$. The bounds of correctness are roughly equivalent to these described in Lemma 1 and Theorem 1. However, the term $(1/4)^{r}$ now takes a dependency on the probability of generating a string that is far from the desired constraint.
C.2 $k$-ary Label Sets
To convert a dataset of arity $k>2$, it is necessary to rerun the No-Data algorithm k1k-1 times. On each run, the label set $\{l_{1},\dots,l_{k}\}$ is mapped as $\{l_{1}\},\{l_{2},\dots,l_{k}\}\mapsto\{0\},\{1\}$. In these cases, the runtime of the algorithm is increased by a factor of k1k-1.
@@ Appendix D Full Results
D.1 Ablation: Lying Evaluators and Lying Generators
The results are in Table 5. Throughout that study DT randomly (w.p. $1/10$) flipped the label when evaluating. The lies tested were (in order in the table) when the model claims to
  1. 1.
  2. 2.
  3. 3.
The last ‘lie’ is more akin to how a realistic generator would behave, sometimes understanding the challenges well, but failing them with some probability.
D.2 Ablation: LLM Generation Strategies
Given the LLMs’ ability to follow instructions, it is likely that their performance could be sensitive to how the datapoints are generated during the EV protocol. Hence in this study we compared the performance of both the LLM whose results we presented in the main results (o3-mini) with another LLM, GPT-4o [OpenAI, 2024]. We tested two different generation strategies: picking a matching datapoint from the data, as in Section 6.1.2, and actually generating a datapoint from the rubric. The results are in Table 3. We found that, again, the experimental results adjusted to the predictions, with lying and low-performing models being easily spotted by the No-Data Algorithm. However, we also found that actual generation led to bad performance in both LLMs. We note that this is expected: in general, classifying an element in a language is much harder than generating it [Kleinberg and Mullainathan, 2024, Kalavasis et al., 2025].
o3-mini (IP) o3-mini (OOP) GPT-4o (IP) GPT-4o (OOP)
Picking
Accuracy/F1 97.6 / 97.6 59.0 / 70.0 58.0 / 67.1 56.4 / 65.4
Successes/Flips 81.3 / 1.8 27.9 / 6.0 28.7 / 71.3 10.8 / 8.8
Generating
Accuracy/F1 94.2 / 94.2 60.6 / 65.5 56.6 / 63.8 53.4 / 63.6
Successes/Flips 55.4 / 5.0 28.7 / 7.8 5.8 / 10.8 34.5 / 4.6
Table 3: Ablation study on LLM generation strategies, comparing a high-performing model (o3-mini) with a lower-performing model (GPT-4o) as a LLM-based baseline. We report the highest-performing run at $\phi=0.1$. Both models were able to broadly adjust to their original accuracy, but success rates are markedly different when the models are generating a datapoint versus picking one.
We also performed a study on the overall performance of various models under the generation strategy, comparing them to the known case. This study includes results with DeepSeek [DeepSeek-AI et al., 2025], GPT-4o, and Qwen 2.5 VL 7B [Qwen Team, 2025] (Table 4) Outside of o3-mini, all models–open or not–struggled with this problem, even in the known case. Consequentially, their performance in the unknown case also adjusted to the predictions. When the accuracy was high in OOP settings, the success rate was low–as expected.
@@ D.2 Ablation: LLM Generation Strategies
Table 4: Comparison of the No-Data algorithm with respect to various LLMs and values of $\phi$. The known case reflects the LLM’s expected performance in the dataset, while the unknown case shows what the No-Data Algorithm outputs. In the unknown case, all models broadly adjusted to their known case. The models also showed on average low accuracies across the board. However, success rates were markedly different between the known and unknown cases under high accuracy, indicating possible deception. When the accuracy was low, success rates remained low regardless of the setting. For all of the above, we used the generating strategy.
@@
D.3 Ablation: Linguistic Competency
In this section we present an expanded version of the results from Section 6.2, including the performance of Qwen 2.5B VL 7B on the same corpus, with GPT-4.1 as a verifier. This model has no indications of being proficient in West Frisian, although the calibration of GPT-4.1 in IP indicates that the verifier is capable of acting as a reliable component. Indeed, in the IP case, where the model was given the correct rubric (i.e., it did not lie), the success rate was still low. In the OOP case (where it was set up to lie), the success rate was lowest. As before, this number remained steady throughout the experimentation. The full results are in Table 6.
@@ D.3 Ablation: Linguistic Competency
Table 6: Performance of the LLM on the natural-language version of the No-Data Algorithm with our ablation study including a model not proficient in the tasks. The success rate for this model was low in both the IP case (the model not lying; just wrong) and the OOP case (the model lying and wrong), although much lower in the latter.
Appendix E Rubrics

### v4

@@ Appendix A Detailed Proofs
In this section we provide more careful treatments of the proofs of Lemma 5.1 (Appendix A.1) and Theorem 5.2 (Appendix A.4).
@@ A.1 Correctness of the EV protocol
We separate our proof in two: the first part concerns itself with completeness, or whether the challenges will provide sufficient information to determine if the evaluator can be lying. The second part deals with robustness under lying, cheating, or lucky scenarios by the evaluator.
A.1.1 Part 1 (Completeness)
@@
Lemma A.1.
Let $x,x^{\prime}\in X$ be two binary strings such that $x\neq x^{\prime}$. To determine if the evaluator is lying, the checks from Challenge 1 and Challenge 2 are necessary and sufficient. In other words, it is necessary and sufficient to check if $x$ is similar to $x^{\prime}$.
@@
(Necessary) It follows from the definition of $C$. Since $f(x)=f(x^{\prime})$, the check for $x\equiv x^{\prime}$ is a necessary condition, but not sufficient. This challenge yields information around the internal structure of $x,x^{\prime}$ up to permutation. It is necessary because it is possible to have two $x,x^{\prime}$ such that $x^{\prime}\not\in\{\pi_{i}(x)\}$ (i.e., two strings that aren’t permutations of one another), and yet $C(x^{\prime})=C(x)$. However, this does not prove that $x$ is similar to $x^{\prime}$, since there could be a subset $t\in S_{x^{\prime}}.t\not\in S_{x}$ and $c(t)=c(s)$ for some $c\in\mathcal{C}$ (for example, if $c$ contains a disjunctive clause). Hence running this challenge provides information as to whether the encoding of $C(x^{\prime})$ contains all information fed into $\sigma$ for a specific $x,x^{\prime}$. It accounts for variation in internal structure of the string so that $C(x)=C(x^{\prime})$, but does not account for the ordering of these bitstrings.
@@
(Sufficient) It follows directly from the above. Knowing $C(x)=C(x^{\prime})$ solely implies that the final encodings for $x$ and $x^{\prime}$ are equivalent, not that they are structurally equivalent. Conversely, knowing $\bar{C}(x),\bar{C}(x)$ does not necessarily yield information as to how these are converted into $C(x)$ and $C(x^{\prime})$. Hence it is sufficient to know both quantities to determine if $x$ is similar to $x^{\prime}$.
Consequentially, testing if $x$ is similar to $x^{\prime}$ naturally implies that these challenges are sufficient to test $f(x)=f(x^{\prime})$ (via $x\equiv x^{\prime}$), and necessary (via $x\cong x^{\prime}$) to determine if the evaluator is lying.
@@ A.1.2 Part 2 (Robustness)
There are then two (disjoint) possibilities for a generated $\langle x^{\prime},y^{\prime}\rangle$: either that the evaluator is lying, but gets lucky; or that it is cheating and can generate the answer for one of the challenges. Remark that generating an answer to pass both challenges is equivalent to knowing $f$. We thus focus on the two possible lies from the evaluator, which would allow it to pass either challenge.
@@ A.3 Lie 2: $f(x^{\prime})\neq y^{\prime}$
It follows from Lemma A.1 that both challenges are necessary and sufficient to decide if an evaluator is lying; and from Lemma A.2 that the ability of the verifier to do this is bounded by a probability of error of $(1/4)^{r}$. Putting both lemmas together we obtain the proof for Lemma 5.1.
@@ A.4 Correctness Bounds of the No-Data Algorithm
For this, it is easier to work with errors $\epsilon=1-\alpha$, and remark that, from Lemma 5.1 the first term (undetected, no flip) is given by
@@ Appendix B Characterisation Bounds
Straightforward by noting the number of calls in the algorithm. ∎
Remark B.2.
The runtime of the No-Data Algorithm is linear (up to a factor of $r$), but this algorithm is designed to calibrate an evaluator, not label a full dataset. Thus, $D$ and $r$ may be small, as in Section 6.2.
@@ Appendix C Extensions
C.1 $k$-ary Label Sets
To convert a dataset of arity $k>2$, it is simply a matter of following a ‘one-vs-all’ breakdown, rerunning the No-Data algorithm k1k-1 times. On each run, the label set $\{l_{1},\dots,l_{k}\}$ is mapped as $\{l_{1}\},\{l_{2},\dots,l_{k}\}\mapsto\{0\},\{1\}$. In these cases, the runtime of the algorithm is increased by a factor of k1k-1, but the theoretical guarantees hold.
It is then worthwhile considering whether it is possible to circumvent the runtime by simply running the No-Data algorithm without any modifications, except the label flip, and still obtain similar results. To test this, we ran the same setup from Section 6.2, but splitting the aggregator into three labels:
  1. 1.
    Label 0: All of the $c_{i}\in\bar{C}$ evaluate to zero.
  2. 2.
    Label 1: All, except one, of the $c_{i}\in\bar{C}$ evaluate to zero.
  3. 3.
    Label 2: None of the $c_{i}\in\bar{C}$ are zero.
The results are in Table 4. Overall, it is possible to see that the results largely adhere to the previously-observed binary label settings. Namely, lying evaluators have near-zero successes and their original performance is close to the known case, with a best-of difference of 6% accuracy. This difference is comparable to the binary case in Table 8 for an untuned $\phi$. Indeed can observe a slight improvement with respect to the known case when the label flip is disabled. As shown in Section 5 (and experimentally in Appendix D.3), the label flip is required to achieve full restoration of the original score, which will require further algorithmic developments for a $k$-ary setting, and are beyond the scope of this work. In sum, while empirically the No-Data Algorithm may be used for $k>2$, further theoretical development is needed to extend this algorithm to non-‘one-vs-all’ strategies with the same guarantees.
We then may pose the following:
Conjecture 1:
Let $Y$ be a label set with arity $k>2$, and $D$ a dataset. Then, for any number of rounds $r$, there exists an $n$ such that a version of the No-Data Algorithm runs in $\text{O}(nr|D|)$ steps, where 1<m<n<k11<m<n<k-1 and $m$ depends on $k$.
For a ‘version’ of the No-Data Algorithm, we mean that it maintains the same correctness bounds from Theorem 5.2.
Table 4: Ablation study reframing the setup from Section 6.2 into a ternary-labelled experiment (0,1,2{0,1,2}). We report the results with accuracy and $F_{\mu}$ (macro) score. For comparison, a random guesser obtains 33 accuracy and 30 $F_{\mu}$. In this experiment, the known case was still reconstructable from the algorithm’s run, although with more noise. Disabling label flips improved restoration of the original scores. Full reconstruction would require the development of further algorithmic strategies to retain all guarantees from the No-Data Algorithm.
IP Acc. / Fμ OOP Acc. / Fμ IP Succ. / Flips OOP Succ. / Flips
Known
Predicting $y$ 71.8 / 61.1 10.5 / 7.0 –/– –/–
Unknown
$\phi=0.1$
   Flip 66.0 / 55.7 12.2 / 10.8 39.4 / 4.3 0.4 / 11.3
   No flip 67.0 / 57.1 10.1 / 6.4 39.4 / – 0.4 / –
$\phi=0.3$
   Flip 59.8 / 52.3 20.0 / 17.8 38.1 / 18.1 0.8 / 34.2
   No flip 61.6 / 55.1 10.5 / 7.0 38.1 / – 0.8 / –
$\phi=0.7$
   Flip 39.5 / 40.0 30.9 / 21.1 35.3 / 47.4 0.8 / 71.6
   No flip 46.8 / 43.4 0.10 / 6.7 35.3 / – 0.8 / –
C.2 Distance-based Constraints
The constraint that $C(x)=C(x^{\prime})$ is too strict for some applications. We sketch out in this section a variant of the No-Data Algorithm where, instead of requiring equality, we request closeness with respect to some arbitrary metric $\delta$ in Challenge 2. Suppose now that the failure is given whenever $\delta(C(x),C(x^{\prime}))\geq\epsilon$, for some $\epsilon\in[0,1/2]$. The bounds of correctness are roughly equivalent to these described in Lemma 5.1 and Theorem 5.2. However, the term $(1/4)^{r}$ now takes a dependency on the probability of generating a string that is far (based on $\delta$) from the desired constraint, and thus the bounds become weaker when said probability is larger than $1/2$.
@@ Appendix D Full Results
D.1 LLM-Based Results
In a full version of our study from Section 6.1, we evaluated multiple open and closed LLMs. The LLMs studied were o3-mini (as reported in the main section), DeepSeek (DeepSeek-AI et al., 2025), GPT-4o, and Qwen 2.5 VL 7B (Qwen Team, 2025), all under the same experimental setup from Section 6.1. The results can be found in Table 5. Outside of o3-mini, all models struggled with this problem, even in the known case. Consequentially, their performance in the unknown case also adjusted to the predictions. It is worth noting that, even when their accuracy was (relatively) high in OOP, their success rate was low–as expected. This is particularly noticeable in o3-mini, GPT-4o, and DeepSeek.
Table 5: Comparison of the No-Data algorithm with respect to various LLMs and values of $\phi$. The known case reflects the LLM’s expected performance in the dataset, while the unknown case shows what the No-Data Algorithm outputs. All LLMs broadly adjusted to their known case. The LLMs also showed on average low accuracies across the board. However, success rates were markedly different between the known and unknown cases under high accuracy, indicating possible deception. When the accuracy was low, success rates remained low regardless of the setting.
@@ D.1 LLM-Based Results
D.2 Ablation: Lying Evaluators and Lying Generators
In this study we evaluated the impact of various aspects of lying in evaluation, both during generation and evaluation (of $C$). The results are in Table 6. The lies tested were when the model claims to:
  1. 1.
  2. 2.
  3. 3.
The last ‘lie’ is more akin to how a realistic generator would behave (e.g., an LLM), sometimes understanding the challenges well, but failing them with some probability.
@@ D.2 Ablation: Lying Evaluators and Lying Generators
D.3 Ablation: LLM Generation Strategies
In this study we accounted for the sensitivity to the prompt. This is because it is likely that an LLM’s performance in the EV protocol could be marked by how it generates the datapoints. We tested two different generation strategies: picking a matching datapoint from the data, as in Section 6.1.2; and actually generating a datapoint from the rubric. The dataset is the synthetic dataset, and we compared the performance of the LLM from Section 6.1 (o3-mini); and GPT-4o (OpenAI, 2024). The results are in Table 7. We found that, again, the experimental results adjusted to the predictions, with lying and low-performing models being easily spotted by the No-Data Algorithm. However, we also found that actual generation (versus picking) led to lower performance in both LLMs; with o3-mini having a slightly lower drop (-3%) than GPT-4o (-5%).
Table 7: Ablation study on LLM generation strategies, comparing a high-performing model (o3-mini) with a lower-performing model (GPT-4o). We report the results at $\phi=0.1$. Both LLMs were able to broadly adjust to their original accuracy, but success rates are different when they are generating a datapoint versus picking it.
o3-mini (IP) o3-mini (OOP) GPT-4o (IP) GPT-4o (OOP)
Picking
Accuracy/F1 97.6 / 97.6 59.0 / 70.0 58.0 / 67.1 56.4 / 65.4
Successes/Flips 81.3 / 1.8 27.9 / 6.0 28.7 / 71.3 10.8 / 8.8
Generating
Accuracy/F1 94.2 / 94.2 60.6 / 65.5 56.6 / 63.8 53.4 / 63.6
Successes/Flips 55.4 / 5.0 28.7 / 7.8 5.8 / 10.8 34.5 / 4.6
D.4 Ablation: Linguistic Competency
In this section we present an expanded version of the results from Section 6.2, including the performance of Qwen 2.5B VL 7B on the same corpus, with GPT-4.1 as a verifier. This model has no indications of being proficient in West Frisian. On the other hand, the calibration of GPT-4.1 in IP through the known case indicates that this verifier is capable of acting as a comparatively reliable component. Indeed, in the IP case, where the model was given the correct rubric (i.e., it did not lie), the success rate for the Qwen-based evaluator was low–thus indicating inability to solve the problem, as opposed to deception. Consequentially, in the OOP case (where it was set up to lie), the success rate was lowest. As before, this number remained steady throughout the experimentation. The full results are in Table 8.
Table 8: Performance of Qwen 2.5B and GPT-4.1 on the natural-language version of the No-Data Algorithm with our ablation study. In this setup, Qwen 2.5B and GPT-4.1 act as evaluators, and GPT-4.1 as a verifier. The success rate for Qwen 2.5B was low in both the IP case (implying that the the model is not lying; just wrong) and the OOP case (where Qwen 2.5B was lying and wrong). Since the success rate is much lower in the latter, it follows that the No-Data Algorithm is particularly proficient at detecting confident-but-wrong evaluators.
@@ D.4 Ablation: Linguistic Competency
D.5 Comparison With Baselines
It was noted in Section 6 that baselines for this problem remain challenging, as they are unable to spot lying evaluators. Ensemble methods, for example, provide ‘a’ label, but not a measure of trust. Typically, this would be given by agreement metrics such as pairwise Cohen’s $\kappa$ or percentage agreement (PA). However, these metrics must be interpreted and calibrated, which in turn depends on the problem at hand (Braylan et al., 2022). Likewise, they have other drawbacks, such as sensitivity to class imbalance (Cicchetti & Feinstein, 1990; Gwet, 2010). These issues are hard to circumvent when there do not exist labels or references for calibration. On the other hand, in this scenario the No-Data Algorithm matches the known-case performance, provides a trust metric (the success rate), and spots deception.
In this section we illustrate this by comparing two consensus-based strategies (ensembles): aggregating the predictions of three evaluators (GPT-4.1, o3-mini, and Qwen-2.5) by averaging them, and by taking a majority vote. The comparisons are with respect to the known case from Section 6.2, for GPT-4.1 and including IP and OOP. The agreement metrics are Cohen’s $\kappa$ and PA. The full results are in Table 9.
Perhaps unsurprisingly, in terms of baselines, ensembling three evaluators matched or outperformed the accuracy and F1 for the (single model) IP known case. Nonetheless, they were marked as presenting poor agreement ($\kappa=0.38$; PA = $0.72$), as per Fleiss & Cohen (1973). Without calibration, poor agreement would indicate caution. Indeed, this value should be on average higher since the classes are balanced and the ensemble performances are good (Sim & Wright, 2005).
Moreover, when in the presence of lying evaluators (OOP), ensembling methods had a difficult time. Averaging their predictions correctly predicted random performance (50.1% accuracy); but majority vote aggregation reported accuracies well above what would be expected (61.6%), thus indicating vulnerability to deception. Even though Cohen’s $\kappa$ was reliable in this case (noting only chance agreement, $\kappa=0.0$), PA reported a moderate agreement ($0.37$). In practice, these contradictory findings would require calibrating the metrics, especially if the IP case’s $\kappa$ is used as justification for the OOP case. This is not possible without labels. [^5] Realistically, in an imbalanced-labels or multi-class scenario, metrics such as PA and Cohen’s $\kappa$ are even more sensitive, thus further complicating calibration (Cicchetti & Feinstein, 1990).
In contrast, the No-Data Algorithm provided a way to accept/reject the evaluator’s predictions (the success rate; 86.2 for IP and 1.2 for OOP) as well as a slightly higher accuracy (76.9%) when compared to the single-model IP case. Since no calibration is needed for any parameter, we argue that, in the unknown case (the focus of our work), it is much more helpful to use this method over other algorithms.
Table 9: Baseline comparisons between two ensemble methods and the results of the No-Data Algorithm from Table 8. For measures of trust, we report Cohen’s $\kappa$ and PA for the ensembles and success rate for the No-Data Algorithm. In the IP case, Cohen’s $\kappa$ marks agreement as poor (as per Fleiss & Cohen 1973); even when the ensemble’s performance nears or surpasses the known case. In OOP, the ensembles obtained above-random performance, which is impossible given the setup. Thus, trusting the value of a calibrated $\kappa$ in IP for OOP would yield spurious results. Hence better calibration is needed, since the information provided by the agreement metrics is insufficient (i.e., it is either contradictory or does not indicate full agreement). This is, however, not possible in the unknown case. In contrast, the No-Data Algorithm does not rely on calibrations and is robust to lying evaluators, while also being able to restore the known case accuracy.
IP trust OOP trust IP Acc. / F1 OOP Acc. / F1
Known
GPT-4.1 76.3 / 80.3 51.5 / 67.2
No-Data Algorithm 86.2 1.2
GPT-4.1 at $\phi=0.1$ 76.9 / 79.8 50.7 / 16.4
Ensembles 0.39 / 0.72 0.0 / 0.37
Average 80.6 / 81.3 50.1 / 0.0
Majority vote 75.5 / 79.9 61.6 / 48.7
Appendix E Detailed Experimental Methods
For our first experiment, we used o3-mini (version 2025-01-31) with maximum tokens set to 50,000 and left all other parameters as default. GPT-4o (version 2024-05-13) and Qwen 2.5 VL 7B were called with maximum tokens set to 1,024 and temperature set to zero. DeepSeek (version r1-distill-qwen-32b) also had the temperature set to zero, but the maximum tokens set to 2,048 to account for the longer reasoning trace. For the decision tree we used the implementation from sklearn [^6] https://scikit-learn.org/stable/index.html with all parameters, except the random seed, left as default.
All LLMs were called via the Azure OpenAI API. All data analysis was performed on a consumer-grade laptop. Random seeds are set for the training in the code and indicated in the repository.
E.1 Rubrics